A = 36m squared, b = , h = 12 m
Found 2 solutions by jim_thompson5910, stanbon:
Answer by jim_thompson5910(35256)

A = (bh)/2

36 = (b*12)/2

36*2 = 12b

72 = 12b

72/12 = b

6 = b

b = 6

So the base is b = 6
--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------

Answer by stanbon(75887)

A = 36m squared, b = , h = 12 m
-----------
Area = (1/2)b*h
36 = (1/2)b*12
36 = 6b
b = 6 meters
--------------------
Cheers,
Stan H.
========